# Muchisim: A Simulation Framework for Design Exploration of Multi-Chip Manycore Systems

### v1

@@ Abstract
Current design-space exploration tools cannot accurately evaluate communication-intensive applications whose execution is data-dependent (e.g., graph analytics and sparse linear algebra) on scale-out manycore systems, due to either lack of scalability or lack of detail in modeling the network. This paper presents Muchisim, a novel parallel simulator designed to address the challenges in exploring the design space of distributed multi-chiplet manycore architectures for communication-intensive applications. We evaluate Muchisim at simulating systems with up to a million interconnected processing elements (PEs) while modeling data movement and communication in a cycle-accurate manner. In addition to performance, Muchisim reports the energy, area, and cost of the simulated system, and it comes with a benchmark application suite and two data visualization tools.
Muchisim supports various parallelization strategies and communication primitives such as task-based parallelization and message passing, making it highly relevant for architectures with software-managed coherence and distributed memory. Via a case study, we show that Muchisim helps users explore the balance between memory and computation units and the constraints related to chiplet integration and inter-chip communication. Muchisim enables scaling up the systems in which new techniques or design parameters are evaluated, opening the gate for further research in this area.
@@ I Introduction
The end of Moore’s law and the increase in dataset sizes to be processed have driven the recent surge in development of scale-out manycore systems [15, 29, 91, 38, 17, 101, 43, 94]. Design-space exploration for these architectures at large scales employs roofline and analytical models to estimate performance, by relying on the deterministic nature of communications as well as computations for many machine learning and dense applications. Alternatively, for data-dependent communication-intensive (DCI) applications, such as graph analytics, sparse linear algebra, and database operations, the communication is not deterministic, requiring simulation tools to accurately estimate performance. However, current simulators cannot evaluate DCI applications at scale due to the limited scalability of model processing units in detail. Because of that, recent work on architecture designs for DCI applications are evaluated on relatively small scales, and thus, their network design—whose performance impact would be exacerbated at larger scales—is not well explored [18, 64, 61, 60, 30, 56, 19, 52].
@@
Current manycore architectures typically emphasize speeding up computation, dedicating significant chip area to it at the expense of memory and network resources. However, when the workload is highly parallelized, data movement starts to become the limiting factor for performance. This is especially true for DCI applications but also other application domains such as spectral methods [65], which necessitate a careful balance between computational, memory and network resources. Thus, in manycore systems with thousands to millions of cores, factors such as the number of cores per silicon die, the number of chiplet dies per package, memory hierarchy (integration and coherence), network-on-chip (NoC) and inter-chip interconnect, become critical design considerations. Lack of ability to consider these at large scale can create significant gaps between designed systems and their actual realizations.
@@
To address these issues, we introduce Muchisim,  [^1] Muchisim is derived from the Spanish word muchisimo (very much), alluding to its ability to simulate a large number of PEs in a multi-chip setting. Muchisim is pronounced as ”moo-chee-sim”. a parallel simulator that allows exploration of the manycore architecture design space, with a special focus on communication-intensive applications. [^2] At large scale, memory-intensive applications lead to a communication bottleneck, so we englobe them in the communication-intensive category. Muchisim scales up to millions of interconnected PEs by relying on user instrumentation for code runtime on the compute units while simulating data movement and communication in a cycle-accurate manner. It also includes energy, area and cost models for multi-chip distributed systems, which are critical to reason about design tradeoffs. In addition, Muchisim includes an application benchmark suite and data visualization tools to facilitate the analysis of its outputs.
@@
As the industry leans towards explicit data movement for large manycores due to the prohibitive cost of hardware-based coherence, Muchisim becomes particularly relevant. Muchisim is applicable for architectures with software-managed coherence and distributed-memory, providing an accurate representation of modern scale-out systems. Muchisim facilitates exploring the integration granularity and balance between memory and computation units, as well as constraints related to chiplet integration and inter-chip communication. With a multifaceted analysis aimed at expanding current understanding of manycore architectures for communication-intensive applications, Muchisim offers a flexible yet powerful framework for exploring many design configurations, thereby contributing to the design and optimization of future large-scale systems.
@@
  • A novel performance modeling approach for distributed manycore architectures that allows for scalable simulations of DCI applications across millions of PEs.
@@
  • A benchmark suite of eight applications and ten datasets of different sizes and characteristics.
  • Two visualization tools to compare certain metrics for different datapoints (DUT configurations, applications and datasets) or analyze them for every PE throughout the execution of a single datapoint.
@@
  • Muchisim is the first open-source simulator that accurately simulates DCI applications with billion-element datasets parallelized across a million PEs within tens of hours.
  • Its parallelization achieves linear speedups up to the thread count equal to the columns of the PE grid being simulated.
  • Muchisim accurately models the runtime and area, as validated with the Cerebras Wafer-Scale Engine.
  • Muchisim is a powerful design-exploration tool that helps identify optimal configurations for different metrics (e.g., performance-per-dollar) and applications
The rest of the paper is organized as follows: section II provides background on existing simulation approaches and motivates our work; section III details the architecture class that Muchisim simulates, its design and capabilities, and the benchmark and visualization tool that it includes; section IV presents the validation and scalability analysis of Muchisim, in addition to a case study that uses Muchisim to study different memory integrations; and section V concludes the paper and discusses future work.
@@ II Background and Motivation
As the number of processing elements (PEs) used to parallelize a workload increases, the network bandwidth and topology become important design considerations. Depending on the architecture, the bottleneck may also shift or aggravate with the level of parallelism. Thus, detailed simulation of large manycore systems is critical for design-space exploration, to avoid making suboptimal design decisions that may be hard to correct later in the design process.
@@
The computer architecture community has seen a lot of promising research on hardware-software co-designs and optimizations for data-dependent communication-intensive (DCI) applications [90, 64, 30, 78, 72, 1, 60, 61, 18, 76, 35, 96, 69, 66, 23]. However, these ideas are often evaluated on a limited number of PEs (hundreds to a few thousand) or relatively small datasets (up to millions of data elements), as increasing either of these dimensions increases the simulation time significantly. With Muchisim, we want to enable the exploration of these ideas on larger scales, as well as offer key parameters to explore the balance of the hardware resources and power allocated to memory, network and compute units. For the rest of this section, we outline the desired features of a scale-out manycore system simulator (and their potential trade-offs) and review existing simulators.
Accuracy is one of the most important features of an architecture simulator, as it dictates the reliability of design experiments. For systems of different scales and target application domains, focusing on modeling some of their components in detail may become necessary when building a simulator that balances simulator performance and accuracy [99].
@@
Full versus application-level simulation: Full-system simulators offer the ability to evaluate the performance of a target design in the context of other system components as well as a complete software stack including the operating system (OS). However, this comes at a great expense of speed and scalability even for small-scale systems. Nevertheless, when the execution significantly relies on OS and I/O processes, full-system simulators are essential when evaluating the performance of a design. Simulators such as GEM5, SimFlex, COTSon, RAMP Gold are some of the most commonly used full-system simulators with varying scope and capabilities [12, 31, 7, 92]. In the manycore architectures we are considering in this study, PEs are not expected to run the OS but rather behave as accelerators to host CPUs, and they do not implement hardware coherence in order to save area and power but instead have an explicit view of the memory. In addition, PEs rely on software to orchestrate the data movement—optimized based on application specifics [15, 69, 70, 21, 94]. This makes an application/user-level simulator for these systems more appropriate. Application-level simulators can scale to higher degrees of parallelism than full-system simulators by simulating up to a couple of thousand PEs [55, 80, 26, 3, 7]. However, these simulators cannot efficiently simulate systems beyond this size, and some of them lack detailed network modeling, which significantly affects the accuracy of larger system simulations. BigSim [102] is a simulator originally developed to simulate Blue Gene that performs detailed network simulation after the emulation of the program on a real system to account for network traffic, achieving scaling in tens to hundreds of thousands of cores. BigSim relies on the program to be deterministic, but certain DCI workloads like graph algorithms do not converge deterministically unless additional synchronization steps are employed.
@@
Simulating DCI applications on manycores: Our goal is to model a class of tiled, distributed manycore architectures. This class of architectures is composed of an interconnected number of machine nodes, each with a board of chip packages, further subdivided into chiplets and PEs. Therefore the entire compute system can be viewed as a hierarchically connected grid of PEs. Muchisim is built as a tool to explore the design space of this class of architectures, with a focus on DCI applications. The challenge with these applications is that their data-dependent execution requires functional simulation to precisely capture performance and energy usage. When parallelizing DCI applications across thousands of PEs and billions of data elements, faithfully simulating all architecture components every cycle becomes infeasible.
@@
PE detail level: Since the aforementioned literature has identified the memory bandwidth and inter-PE communication to be the main bottlenecks of these applications, Muchisim focuses on modeling the network and memory systems, while executing the task on the host and relying on the user to provide a performance model for the compute time (e.g., instructions per cycle). A similar approach has been previously used by PriME [26], but PriME simplifies the PE model by simulating each instruction as 1 cycle, whereas Muchisim can use that or a detailed user-provided model (section III). Because the PEs used in scale-out manycores do not interact with the OS and have a simple view of the memory system, it is possible to model the timing of execution relatively accurately using user-provided performance models and without detailed core simulations. Simulators such as SimpleScalar [8], SimFlex [31], and Gem5 [12] offer detailed core modeling with a wide range of supported ISAs. In addition, simulators such as MosaicSim [53] and SimTRaX [84] provide LLVM-based instruction sets. When precise PE models are desired, these simulators can be used to provide the PE- and task-specific cycle counts to be incorporated into Muchisim simulations.
@@
Detailed Network simulation: Since inter-PE communication becomes a significant performance factor for large-scale manycore systems, the capability to model different configurations of the network-on-chip (NoC) and the inter-chip interconnect in detail is essential when modeling these systems. While there are several cycle-accurate NoC simulators such as Xpipes [10], NOXIM [14], SICOSYS [77], BookSim [36], among others [32], most of the existing multicore system simulators [87, 26, 55] do not model the NoC in detail. Although some implementations incorporate these NoC models within system simulators (e.g. Xpipes with Gem5), a functional architecture simulator (with cycle-accurate NoC modeling) able to simulate beyond a thousand PEs was not offered so far.
@@
Scope: Muchisim comes to fill this gap; the class of architectures for which Muchisim is applicable includes—but not limited to—tiled accelerators [69, 68, 103, 4, 18] and dataflow machines [15, 21, 29, 94, 25]. Muchisim can also be used to simulate systolic array architectures like Google’s TPU [38], but these designs usually execute deterministic workloads with near-neighbor communication, and thus they would be modeled much faster using analytical models. Unlike general-purpose simulators for HPC such as SST [79], focusing on this architectures class allows Muchisim to be a highly-scalable, parallel, light-weight simulator while providing detailed simulations of critical aspects of DCI workloads such as NoC modeling at flit-level granularity. Muchisim could also potentially be used to model the performance of workloads running on more general-purpose manycores [101, 50, 27, 20] and server-class chip [22, 5, 58], provided that they use explicit data movement or message passing instead of shared-memory across PEs.
@@ III Muchisim Simulation Framework
In section III-A we first present the hardware components of Muchisim’s target architecture class, and how they are hierarchically organized. Then, section III-B explains the execution models supported in Muchisim and how to describe applications. Next, section III-C elaborates on how Muchisim simulates the execution of the applications on the target architecture. Muchisim is a parallel simulator written in C/C++. Muchisim does not use external libraries beyond threading, so it can be compiled with any C++ compiler that includes OpenMP (-fopenmp) or pthread support (-lpthread). The code can be found in our repository[67] under the src folder, where the configs subfolder contains the files to configure the DUT to simulate, and its latency, energy, area and cost parameters (described in section III-D and section III-E). The gui and plot folders contain the visualization tools to analyze and compare simulation logs (section III-F). The apps and datasets folders contain the benchmark suite of applications and datasets (section III-G). In addition, the repository includes a Readme file explaining how to configure and run Muchisim for different experiments.
@@ III-A The Target Architecture Class
Figure 1: Hierarchical overview of how tiles can be organized on the Muchisim’s target architecture. The board of a cluster node may contain one or multiple chip packages, each with one or multiple chiplets. Packages can be composed of only compute chiplets (with a grid of tiles), or also DRAM chiplets (adjacent to the compute chiplets). Chiplets also include the PHY for inter-chiplet communication. A tile contains one or more processing units (PUs), a network router (R), the task scheduling unit (TSU), and an SRAM scratchpad (SPM). The task queues are mapped into the SPM.
Throughout the paper, we refer to the simulated architecture as the design under test (DUT). We borrow this acronym from the hardware verification community, where it is used to refer to the design that is being tested. We call the machine that runs the simulator, the host, and use DUT’s host to refer to the machine for which the DUT may behave as an accelerator.
@@
Processing Units (PUs): A PU can be an ISA-programmable core, a CGRA unit, or a hardware accelerator, depending on the performance, energy and area model provided. We have only evaluated cases where the PUs are homogeneous across the chiplet, but Muchisim could potentially simulate heterogeneous manycores [27, 21, 16], since a different performance model could be provided for different PU IDs.
@@
Network Routers: Routers generally have five bidirectional ports: North, South, East, West, and the PU ports (connected to the PU’s input and output queues). When configuring hierarchical inter-chiplet networks [66] or ruche channels [71, 39], some or all routers have an additional set of cardinal ports (for a total of nine). Muchisim supports 2D mesh and 2D folded torus topologies[24] (with dimension-ordered routing) across tiles and chiplets, up to the level of a node. Across nodes, Muchisim currently supports a mesh interconnect, but it could be extended to support other cluster-level topologies [11, 47, 41, 9, 2]. In addition, Muchisim supports configuring reduction trees of different degrees [70] .
@@
Task Scheduling Unit (TSU): To support task-based parallelizations and message-triggered tasks (see section III-B), Muchisim considers a TSU per tile, which takes messages from the input queues (one per task type) and schedules them to a PU. The TSU decides which task to schedule next—from the ones available at the queues—based on a scheduling policy. The policies currently supported are round-robin, priority-based (e.g., prioritizing tasks from a particular queue), or occupancy-based (to prevent full queues from backpressuring the network).
@@
SRAM scratchpad memories (SPMs): The size of each tile’s memory is specified in Kibibytes (KiB); it is the same for all the tiles. Muchisim supports using the SPM as a write-back cache (direct-mapped or associative) when DRAM is integrated on-package or as a scratchpad (given the tile’s local address space) when the chip’s main memory is the tile-distributed SRAM. In cache mode, we model the area and energy of the tags (and valid/dirty bits) using part of the SPM; the cacheline width equals the bitline width of the DRAM memory controller (512 by default). Cache misses go to the on-chip memory controller, which fetches the full cacheline from DRAM directly since there is no hardware coherence in our target DUT. Dirty lines are written back to DRAM upon eviction.
@@
Queues are currently modeled as circular FIFOs mapped into the SPM, which allows for compile-time configuration on the DUT. Thus, in our energy model, queue reads and writes are similar to load and store operations. Alternatively, one could change the model to consider them as separate structures, with the corresponding tradeoff in area and energy.
@@
Prefetching: Muchisim supports modeling next-line and pointer-indirection data prefetching from DRAM into the SPM (used as a cache). The latter is enabled by the fact that tasks can be split at pointer indirection and the TSU can prefetch the data that tasks (waiting in the input queue) will use.
@@
Interconnect links: Muchisim offers parameters to set the width or the number of interconnect links across chiplets or chip packages. Across cluster nodes, one can set the multiplexing factor of the links between nodes, i.e., how many edge tiles share an inter-node link.

### v2

@@ Abstract
The design space exploration of scaled-out manycores for communication-intensive applications (e.g., graph analytics and sparse linear algebra) is hampered due to either lack of scalability or accuracy of existing frameworks at simulating data-dependent execution patterns. This paper presents MuchiSim, a novel parallel simulator designed to address these challenges when exploring the design space of distributed multi-chiplet manycore architectures. We evaluate MuchiSim at simulating systems with up to a million interconnected processing units (PUs) while modeling data movement and communication cycle by cycle. In addition to performance, MuchiSim reports the energy, area, and cost of the simulated system. It also comes with a benchmark application suite and two data visualization tools.
MuchiSim supports various parallelization strategies and communication primitives such as task-based parallelization and message passing, making it highly relevant for architectures with software-managed coherence and distributed memory. Via a case study, we show that MuchiSim helps users explore the balance between memory and computation units and the constraints related to chiplet integration and inter-chip communication. MuchiSim enables evaluating new techniques or design parameters for systems at scales that are more realistic for modern parallel systems, opening the gate for further research in this area.
Index Terms:
graph, sparse, communication-intensive kernels, simulator, design-exploration, multi-chip, scale-out, manycore
@@ I Introduction
The end of Moore’s law and the increase in application dataset sizes have driven the recent surge in development of scale-out manycore systems [16, 2, 92, 38, 18, 102, 43, 95]. Design-space exploration for these architectures at large scales employs roofline and analytical models to estimate performance, by relying on the deterministic nature of communications as well as computations for many ML and dense applications [52]. However, communication pattern in data-dependent communication-intensive (DCI) applications, such as graph analytics, sparse linear algebra, and database operations, is not deterministic, and therefore estimation of performance for running these applications on a system under design requires simulation tools. Current simulators cannot evaluate DCI applications at scale due to the limited scalability of detailed processing unit models. Thus, recent work on architecture designs for DCI applications are evaluated on relatively small scales, and their network design—whose performance impact would be exacerbated at larger scales—is not well explored [19, 65, 62, 61, 30, 57, 20, 53].
@@
Current manycore architectures typically emphasize accelerating computation, dedicating significant chip area to it, at the expense of memory and network resources. However, when the workload is highly parallelized, data movement increasingly becomes the limiting factor for performance. This is especially important when using such manycore systems to accelerate communication-intensive applications such as graph analytics, sparse linear algebra, and spectral methods [70, 66], which necessitate a careful balance between computational, memory and network resources. Thus, in manycore systems with thousands to millions of cores, factors such as the number of cores per silicon die, the number of dies per package, memory hierarchy, network-on-chip (NoC) and inter-chip interconnect, become critical design considerations. The inability to evaluate these aspects at a large scale creates significant gaps between designed systems and their actual realizations.
@@
To address these issues, we introduce MuchiSim  [^1] MuchiSim is derived from the Spanish word muchisimo (very much), alluding to its ability to simulate a large number of PUs in a multi-chip setting. MuchiSim is pronounced as ”moo-chee-sim”., a parallel simulator for exploring the manycore architecture design space, with a special focus on communication-intensive applications [^2] At large scale, memory-intensive applications lead to a communication bottleneck; we call this category of applications communication-intensive.. MuchiSim scales up to millions of interconnected PUs by relying on user instrumentation for code runtime on the compute units while simulating data movement and communication every cycle. It also includes energy, area and cost models for multi-chip distributed systems; these are critical to reason about design tradeoffs. In addition, MuchiSim includes an application benchmark suite and data visualization tools to facilitate the analysis of its outputs.
@@
As industry trends towards explicit data movement for large manycores due to the prohibitive cost of hardware-based coherence [16, 2, 92, 102], MuchiSim becomes particularly relevant. MuchiSim is applicable for architectures with distributed memory and software-managed coherence, which is becoming increasingly popular in modern scale-out systems. MuchiSim facilitates exploring the integration granularity and balance between memory and computation units, as well as constraints related to chiplet integration and inter-chip communication.
With a multifaceted analysis aimed at expanding the current understanding of manycore architectures for DCI applications, MuchiSim offers a flexible yet powerful framework for exploring many design configurations, thereby contributing to the design and optimization of future large-scale systems.
@@
  • A novel performance modeling approach for distributed manycore architectures that allows for scalable simulations of data-dependent and communication-intensive applications across millions of PUs.
@@
  • A benchmark suite of eight applications especially programmed for distributed scale-out systems.
  • Visualization tools that allow comparing system-wide metrics for different evaluations (i.e., design configurations, applications and datasets) and analyzing per-PU metrics throughout the entire execution of a particular evaluation.
@@
  • MuchiSim is the first open-source framework that precisely simulates DCI applications with billion-element datasets parallelized across a million PUs within tens of hours.
  • Its parallelization achieves linear speedups up to a host thread count equal to the number of columns of the manycore grid being simulated.
  • MuchiSim closely matches runtime and area of the real runs of the Cerebras Wafer-Scale Engine when using their reported workload implementation and network specification.
  • MuchiSim is a powerful design-exploration tool that helps identify optimal configurations for different metrics (e.g., performance-per-dollar) and applications.
The rest of the paper is organized as follows: section II provides background on existing simulation approaches and motivates our work; section III details the architecture class that MuchiSim simulates, the design of MuchiSim and capabilities, and the benchmark and visualization tool that it includes; section IV presents the validation and scalability analysis of MuchiSim, in addition to a case study that uses MuchiSim to study different memory integrations; and section V concludes the paper and discusses future work.
@@ II Background and Motivation
As the number of processing units (PUs) used to parallelize a workload increases, the network bandwidth and topology become critical factors in the performance of the system. Depending on the architecture, bottlenecks may also shift or worsen with the level of parallelism. Thus, detailed simulation of large manycore systems is crucial for design-space exploration, to avoid making suboptimal design decisions that may be hard to correct later in the design process.
@@
The computer architecture community has seen a lot of promising research on hardware-software co-designs and optimizations for data-dependent communication-intensive (DCI) applications [91, 65, 30, 79, 73, 1, 61, 62, 19, 77, 35, 97, 70, 67, 24]. However, these ideas are often evaluated on a limited number of PUs (hundreds to a few thousand) or relatively small datasets (up to millions of data elements), as increasing either of these dimensions significantly increases simulation time. MuchiSim enables the exploration of architectural ideas on larger scales, as well as offers key parameters to explore the balance of hardware resources and power allocated to memory, network and compute.
Maximizing accuracy as well as performance is the main objective for the design and engineering an architecture simulator, as it dictates reliability as well practicality of design experiments [100]. For the rest of this section, we outline the critical features of a scale-out manycore system simulator (and their potential trade-offs) and review existing simulators.
II-A Full-system vs. Application-level Simulation
Full-system simulators offer the ability to evaluate the performance of a target design in the context of other system components as well as a complete software stack including the operating system (OS). However, this comes at a great expense of speed and scalability even for small-scale systems. Nevertheless, when the execution significantly relies on OS and I/O processes, full-system simulators are essential when evaluating the performance of a design. Simulators such as GEM5, SimFlex, COTSon, RAMP Gold are some of the most commonly used full-system simulators with varying scope and capabilities [13, 31, 8, 93].
In the manycore architectures we are considering in this study, PUs are not expected to run the OS but rather behave as accelerator processing elements. These manycores save silicon area and power by not implementing hardware coherence and instead having an explicit view of the memory. In addition, PUs rely on software to orchestrate the data movement—optimized based on application specifics [16, 70, 71, 22, 95]. This makes an application/user-level simulator for these systems more appropriate.
@@
Application-level simulators can scale to higher degrees of parallelism than full-system simulators by simulating up to a couple of thousand PUs [56, 81, 27, 4, 8]. However, existing simulators cannot efficiently simulate systems beyond this size, and some of them lack detailed network modeling, which significantly affects the accuracy of larger system simulations. BigSim [103] is a simulator originally developed to simulate Blue Gene that performs detailed network simulation after the emulation of the program on a real system to account for network traffic, achieving scaling in tens to hundreds of thousands of cores. BigSim relies on the program to be deterministic, but certain DCI workloads like graph algorithms do not converge deterministically unless additional synchronization steps are employed.
II-B Simulating DCI Applications on Manycores
This paper focuses on a class of tiled, distributed manycore architectures. This class of architectures is composed of an interconnected number of machine nodes, each with a board of chip packages, further subdivided into chiplets and processing tiles. Therefore the entire compute system can be viewed as a hierarchically connected grid of tiles.
@@ II-B Simulating DCI Applications on Manycores
MuchiSim is a framework to explore the design space of this class of architectures, with a focus on DCI applications. The challenge with these applications is that their data-dependent execution requires functional simulation to precisely capture performance and energy usage. When parallelizing DCI applications across thousands of PUs and billions of data elements, faithfully simulating all architecture components every cycle becomes infeasible.
@@
PU modeling: Since prior literature has identified the memory bandwidth and inter-PU communication to be the main bottlenecks of these applications, MuchiSim focuses on modeling the network and memory systems, while executing compute tasks natively on the host and relying on the user to provide a performance model for the compute time (e.g., cycles per basic block). A similar approach has been previously used by PriME [27], but PriME simplifies the PU model by simulating each instruction as 1 cycle, whereas MuchiSim can use that or a detailed user-provided model (section III). Because the PUs used in scale-out manycores do not interact with the OS and have a simple view of the memory system, it is possible to model the execution using user-provided cycle counts and without detailed core simulations. Simulators such as SimpleScalar [9], SimFlex [31], and Gem5 [13] offer detailed core modeling with a wide range of supported ISAs. In addition, simulators such as MosaicSim [54] and SimTRaX [85] provide LLVM-based instruction sets. When precise PU models are desired, these simulators can be used to provide the task-specific cycle counts to be incorporated into MuchiSim simulations.
@@
Network modeling: Since inter-PU communication becomes a significant performance factor for large-scale manycore systems, the capability to model different configurations of the network-on-chip (NoC) and the inter-chip interconnect in detail is essential when modeling these systems. While there are several cycle-accurate NoC simulators such as Xpipes [11], NOXIM [15], SICOSYS [78], BookSim [36], among others [32], most of the existing multicore system simulators [88, 27, 56] do not model the NoC in detail. Although some implementations incorporate these NoC models within system simulators (e.g. Xpipes with Gem5), a functional architecture simulator (with cycle-accurate NoC modeling) able to simulate beyond a thousand PUs has not been offered so far. MuchiSim fills this gap.
II-C Scope of Applicability
@@
The class of architectures for which MuchiSim is applicable includes—but not limited to—tiled accelerators [70, 69, 104, 5, 19] and dataflow machines [16, 22, 2, 95, 26]. MuchiSim can also be used to simulate systolic array architectures like Google’s TPU [38], but these designs usually execute deterministic workloads with near-neighbor communication, and thus they would be modeled much faster using analytical models.
Unlike general-purpose simulators for HPC such as SST [80], focusing on this architecture class allows MuchiSim to be a highly-scalable, parallel, light-weight simulator while providing detailed simulations of critical aspects of DCI workloads such as NoC modeling at flit-level granularity. MuchiSim could also potentially be used to model the performance of workloads running on more general-purpose manycores [102, 50, 28, 21] and server-class chip [23, 6, 59], provided that the application is written in a way that does not require hardware coherence.
@@ III MuchiSim Simulation Framework
In section III-A we first present the hardware components of MuchiSim’s target architecture class, and how they are hierarchically organized. Then, section III-B explains the execution models supported in MuchiSim and how to describe applications. Next, section III-C elaborates on how MuchiSim simulates the execution of the applications on the target architecture.
MuchiSim is a parallel simulator written in C/C++ which does not use external libraries beyond threading, and thus, it can be compiled with any C++ compiler that includes OpenMP (-fopenmp) or pthreads support (-lpthread). The code can be found in our repository[68] under the src folder, where the configs subfolder contains the files to configure the system to simulate, and its latency, energy, area and cost parameters (described in section III-D and section III-E). The gui and plot folders contain the visualization tools to analyze and compare simulation logs (section III-F). The apps and datasets folders contain the benchmark suite of applications and datasets (section III-G). In addition, the repository includes a Readme file explaining how to configure and run MuchiSim for different experiments.
@@ III-A The Target Architecture Class
Figure 1: Hierarchical overview of how tiles can be organized on MuchiSim’s target architecture. The board of a cluster node may contain one or multiple chip packages, each with one or multiple chiplets. Packages can be composed of only compute chiplets (with a grid of tiles), or also DRAM chiplets (adjacent to the compute chiplets). Chiplets also include the physical layer (PHY) for inter-chiplet communication. A tile contains one or more processing units (PUs), a private local memory (PLM), a network router (R), and a task scheduling unit (TSU). The task queues are mapped into the PLM.
Throughout the paper, we refer to the simulated architecture as the design under test (DUT). We call the machine that runs the simulator, the host. We use DUT’s host to refer to the machine for which the DUT may behave as an accelerator.
@@
Processing Units (PUs): A PU can be an ISA-programmable core, a CGRA unit, or a hardware accelerator, depending on the performance, energy and area models specified. We have only evaluated cases where the PUs are homogeneous across the chiplet, but MuchiSim could potentially simulate heterogeneous manycores [28, 22, 17] since a different performance model could be provided for different PU IDs.
Network-on-chip (NoC): MuchiSim supports evaluating 2D mesh and folded torus topologies[25] (with dimension-ordered routing) for the NoC that connects the routers of every tile. This NoC can span multiple chiplets and chip packages, up to the level of a cluster node. Several physical NoCs can be evaluated, with the same or different NoC widths. Additional NoCs can be regular (i.e., connecting all routers) or Ruche [72, 39] (i.e., connecting to one for every $R$ routers).
@@
Routers generally have five bidirectional ports: North, South, East, West, and the PU ports (connected to the PU’s input and output queues). When configuring hierarchical inter-chiplet [67] or Ruche NoCs, some or all routers have an additional set of cardinal ports (for a total of nine). MuchiSim can also evaluate the Tascade router support for asynchronous and opportunistic reduction trees [71].
@@
Task Scheduling Unit (TSU): To support task-based parallelizations and message-triggered tasks (see section III-B), MuchiSim considers a TSU per tile, which takes messages from the input queues (one per task type) and schedules them to a PU. The TSU decides which task to schedule next—from the ones available at the queues—based on a scheduling policy. The policies currently supported are round-robin, priority-based (i.e., prioritizing tasks from a particular queue), or occupancy-based (to prevent full queues from backpressuring the network).
@@
Private Local Memory (PLM): The size of each tile’s memory is specified in Kibibytes (KiB); it is the same for all the tiles. MuchiSim supports using the PLM as a write-back cache, when DRAM is integrated on-package, or as a scratchpad (given the tile’s local address space) when the chip’s main memory is the tile-distributed SRAM. In cache mode, we model the area and energy of the tags (and valid/dirty bits) using part of the local SRAM; the width of a cacheline is equal to that of the bitline of the DRAM memory controller (set to 512 by default). Cache misses are directed to the on-chip memory controller, which fetches the full cacheline from DRAM directly since there is no hardware coherence in our target DUT. Dirty lines are written back to DRAM upon eviction.
@@
Queues: Inputs and outputs for message triggered tasks (see III-B for details) are mapped into the PLM of each tile and currently modeled as circular FIFOs, which allows for compile-time configuration on the DUT. Thus, in our energy model, queue reads and writes are similar to load and store operations. Alternatively, one could change the model to consider them as separate structures, with the corresponding tradeoff in area and energy.
@@
Prefetching: When DRAM is present in the design, MuchiSim supports modeling next-line and pointer-indirection data prefetching from DRAM into the PLM (used as a cache). The latter is enabled by the fact that tasks can be split at pointer indirection and the TSU can prefetch the data that tasks (waiting in the input queue) will use.
@@
Interconnect links: MuchiSim offers parameters to set the width and the number of interconnect links across chiplets (within a package) and across chip packages. Across cluster nodes, one can also set the multiplexing factor of the links between nodes, i.e., how many edge tiles share an inter-node link. To interconnect nodes, MuchiSim currently supports evaluating a mesh topology, but this could be extended to support other cluster-level topologies [12, 47, 41, 10, 3].